OpenStudy (anonymous):

find Laplace Inverse of : 1/(s^2+6s+13)^2

OpenStudy (anonymous):

Use the fact that L^-1(b/(s+a)^2 + b^2) = e^(-at)sin(bt) Now it's been squared, use convolution integral L(f)L(g) = L(f*g) where f = e^(-at)sin(bt) & g= e^(-at)sin(bt) I know there's a better method, so i'll come back...

OpenStudy (experimentx):

i know a lot better method ... it's called wolfram :P http://www.wolframalpha.com/input/?i=laplace+inverse+1%2F%28s^2%2B6s%2B13%29^2+ ( though currently not working )

OpenStudy (anonymous):

Okay this shuld help... Find laplace transform of sinbt, cosbt, tsinbt and tcosbt... and then u have to look for combinations which will give you the desired result... If u need the full solution, pls tell me. @experimentX : Lolz

OpenStudy (experimentx):

inverse transform are sure a hell of trouble ... :D